A hollow spherical body of inner and outer radii 6 cm and 8 cm respectively floats half-submerged in water. Find the density of the material of the sphere.

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. . Here, Mg = Upward thrust ⇒V pg = (V/2) (pw) * g (where pw = density of water) ⇒ (4/3 πr23 - 4/3 πr13)p = (1/2)( 4/3 πr23)X pw ⇒ (r23 – r13) X p = 1/2r23 X 1 = 865 kg/m3